Reports 1-1 of 1 Clear search Modify search
Detector Operation (Operations Report)
amagazzu - 23:41 Saturday 10 February 2024 (63230) Print this report
Operator Report

The shift was dedicated to the continuation of the planned TCS activity, DAS and CH tuning, carried out by Nardecchia from remote.
At 17:21 UTC we manually unlock the ITF from LOW_NOISE_3. To check and tune the TCS, we locked and unlocked the CITF multiple times before attempting to reach LOW_NOISE_3 again.
ITF locked again from 20:51 UTC to 21:51 UTC, we manually unlocked to verify the locking acquisition. ITF locked again at 22:28 UTC.
As requested by the Commissioning coordinator, at the end of the shift I performed a step in the Etalon WI, changing the setpoint of -0.3 with a ramp of 24 hours (22:36 UTC, command WI Set point [WI_RH_SET=19.78,rampTime=86400]' sent to LSC_Etalon_Acl).
Activity concluded at 22:40 UTC, ITF left locked with AUTORELOCK_FAILSAFE engaged and IN DQSTUDIES.

From 16:52 UTC to 17:19 UTC - CALIBRATED_DF_DAILY

Images attached to this report
Search Help